OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x x = 10 x = 4 x = −14 x = −17 Any help please??

OpenStudy (anonymous):

x=4

OpenStudy (nincompoop):

\[ \frac{x+20}{2} = 3 \] let me teach you how it is done but with another problem \[ \frac{x+35}{5} = 4 \\ 5\times (\frac{x+35}{5}) = 4 \times 5 \\ 1 \times (x+35) = 20 \\ x+35 = 20 x + 35 -35 = 20 - 35 \\x = -15 \]

OpenStudy (nincompoop):

\[ \color {red} 5 \times (\frac{x+35}{\color{red} 5}) = 4 \times \color{red} 5 \\ x+35 = 20 \\ x + 35 \color {red} {- 35} = 20 \color {red} {-35} \\ \sf \large \color {blue} {x = -15} \]